I have trouble answering this simple interest question. A loan company advertises that you only pay 8 cents per day for each $100 borrowed. If a person borrows $500 for 90 days, what amount must the person repay and what annual simple interest rate is the loan company actually charging?(use a 360-day year)
Answer by stanbon(75887)

A loan company advertises that you only pay 8 cents per day for each $100 borrowed. If a person borrows $500 for 90 days, what amount must the person repay and what annual simple interest rate is the loan company actually charging?(use a 360-day year)
------
8 cents per day implies 8*360 = $28.80 per year
Actual simple interest rate = 28.80/100 = 28.8%
-----------------------------
If borrow 500 for 90 days you pay 5*(8 cents)*90 = $36.00
============================================================
Using the formula I = Prt you get:
I = 500*0.288(90/360)
I = 500*0.288(1/4)
I = 500*0.072
I = $36.00
Cheers,
Stan H.